# Attempting an accurate age estimate of the open cluster NGC 6633 using CoRoT and Gaia

### v1

@@ 2 Identifying targets and their properties
Table 1: Information on giant members of NGC 6633
ID HD 170053 HD 170174 HD 170231 HD 170292 TIC 319770092
Symbol in plots cross square circle triangle diamond
Gaia DR3 ID 447746039- 447722337- 447727326- 447724911- 447725330-
1998886144 8527434880 8868842752 3972647168 5860760960
+0700329 +0636003 +0654514 +0642293 +0646000
Dec (degrees) 7.009091 6.600086 6.914277 6.708119 6.766667
[…]
$\nu$_eff 1.4085894 1.4590598 1.4565996 1.4639846 1.4669659
ecl_lat (deg) 30.263 29.84 30.158 29.947 30.007
RUWE 0.962 0.958 1.013 0.879 1.035
Parallax corr. (mas) -0.0329 -0.0328 -0.0326 -0.0325 -0.0324
RV ($\rm km\,s^{-1}$) -29.27(13) -28.71(13) -28.45(13) -29.02(14) -28.68(16)
$G$ (mag) 6.840 7.987 8.328 8.464 8.641
$K_{S}$ (mag)aaaaaaSkrutskie et al. (2006). 4.087 5.674 6.065 6.246 6.443
$G-K_{S}$ (mag) 2.753 2.313 2.263 2.218 2.198
$G_{\rm BP}-G_{\rm RP}$ 1.567 1.295 1.293 1.259 1.242
E(BV)E(B-V) Green bbbbbbFrom Bayestar2019 E(gr)E(g-r), assuming the conversion factor 0.884 given in Green et al. (2019), $A_{G}=2.74\times E(B-V)$ (Casagrande & VandenBerg, 2018), $A_{K_{S}}=0.366\times E(B-V)$, $A_{G_{\rm BP}}=3.374\times E(B-V)$ and $A_{G_{\rm RP}}=2.035\times E(B-V)$ (Casagrande & VandenBerg, 2014). 0.088 0.159 0.150 0.133 0.150
E(BV)E(B-V) Peña ccccccPeña et al. (2017). 0.182 0.182 0.182 0.182 0.182
E(BV)E(B-V) calc. $(V-K_{S})$ddddddCalculated reddening assuming the adopted effective temperatures and requiring agreement with photometric temperatures using Casagrande & VandenBerg (2018). The one using $V-K_{S}$ is the adopted one. 0.048 0.180 0.184 0.177 0.180
E(BV)E(B-V) calc. $(G_{\rm BP}-G_{\rm RP})$ddddddCalculated reddening assuming the adopted effective temperatures and requiring agreement with photometric temperatures using Casagrande & VandenBerg (2018). The one using $V-K_{S}$ is the adopted one. 0.112 0.185 0.205 0.189 0.187
$T_{\rm eff}$ (K)TsTsTsTsTsTsTsantaki et al. (2023). - 4924 5015 5068 5093
$T_{\rm eff}$ (K)MaMaMaMaMaMaMagrini et al. (2021). - 4957 5025 5056 -
$T_{\rm eff}$ (K)CaCaCaCaCaCaCasamiquela et al. (2021), the numbers provided on request by L. Casamiquela (private comm.) - 5010 5095 5135 5165
$T_{\rm eff}$ (K)AdAdAdAdAdAdAdopted effective temperatures. See text for details. 4290 $50$ 5125 5160 5195
$T_{\rm eff}$ from $G_{\rm BP}-G_{\rm RP}$ (K)eeeeeeCalculated using Casagrande & VandenBerg (2018) and assuming the E(BV)E(B-V) from the 3D reddening map of Green et al. (2019). [Fe/H] =+0.0=+0.0 and loggg = 2.7 was assumed. A change of $\pm 0.1$ dex in [Fe/H] corresponds to $\mp$8 K, a change of $\pm 0.5$ in loggg yields $\mp 2$ K 4240 4953 4933 4970 5110
$T_{\rm eff}$ from $G-K_{S}$ (K)eeeeeeCalculated using Casagrande & VandenBerg (2018) and assuming the E(BV)E(B-V) from the 3D reddening map of Green et al. (2019). [Fe/H] =+0.0=+0.0 and loggg = 2.7 was assumed. A change of $\pm 0.1$ dex in [Fe/H] corresponds to $\mp$8 K, a change of $\pm 0.5$ in loggg yields $\mp 2$ K 4368 4987 5023 5029 5104
BCG (mag)ffffffAssuming the adopted temperatures, corresponding reddenings, Casagrande & VandenBerg (2018), and the corrected Gaia Parallaxes. -0.2603 0.0045 0.0194 0.0259 0.0321
[…]
$L(L_{\odot})$ from $K_{S}$ffffffAssuming the adopted temperatures, corresponding reddenings, Casagrande & VandenBerg (2018), and the corrected Gaia Parallaxes. 321.3(53) 122.4(21) 84.7(13) 75.0(12) 63.35(93)
$R_{SB}(R_{\odot})$ffffffAssuming the adopted temperatures, corresponding reddenings, Casagrande & VandenBerg (2018), and the corrected Gaia Parallaxes. 32.5(13) 14.46(48) 11.68(38) 10.85(35) 9.83(32)
A(Li) NLTEMoMaTsMoMaTsMoMaTsMoMaTsMoMaTsMoMaTsA(Li) NLTE from the online Table A5 of Morel et al. (2014) who used NLTE corrections from Lind et al. (2009). For the measurements marked with a *, the abundance was calculated using the numbers of Magrini et al. (2021) or Tsantaki et al. (2023) and the mean difference for the stars in common. 1.44(12) 0.80(12) 1.49(12) 1.21* 1.42*
A(Li) LTEMaMoTsMaMoTsMaMoTsMaMoTsMaMoTsMaMoTsA(Li) LTE from Magrini et al. (2021). For the measurements marked with a *, the abundance was calculated using the number of Morel et al. (2014); Lagarde et al. (2015) or Tsantaki et al. (2023) and the mean difference for the stars in common. 1.17* 0.66(5) 1.31(3) 1.05(3) 1.28*
A(Li) NLTEMaMoWaTsMaMoWaTsMaMoWaTsMaMoWaTsMaMoWaTsMaMoWaTsA(Li) 3D NLTE caluclated from Magrini et al. (2021) LTE values using 3D NLTE corrections from Wang et al. (2021). For the measurements marked with a *, the 3D NLTE correction was applied to the LTE value of Morel et al. (2014); Lagarde et al. (2015) or Tsantaki et al. (2023) corrected to Magrini et al. (2021) using the mean difference for the stars in common. We adopt the larger of the error-bars from Morel et al. (2014) that includes also systematic errors to account for the small general offsets between the samples. 0.97(12)* 0.61(12) 1.26(12) 1.00(12) 1.23(12)*
111 \$MoLaMa\$\$MoLaMa\$footnotetext: A(Li) LTE from the online Table A5 of Morel et al. (2014). For the measurement marked with a *, the abundance was calculated using the number of Magrini et al. (2021) (in the line below) and the mean difference for the two stars in common. \$Ts\$\$Ts\$footnotetext: A(Li) LTE from Tsantaki et al. (2023)
@@ 8 Summary, conclusions, and outlook
The large pixels will also reduce the quality of detailed asteroseismic studies in the crowded fields of clusters from the upcoming PLATO mission (Rauer et al., 2014), which is also not planned to observe the field of NGC 6633. Instead, future missions like STEP444 https://space.au.dk/the-space-research-hub/step/ and Haydn (Miglio et al., 2021b) will allow an extension of asteroseismic studies of cluster stars to a larger scale to ensure proper stellar ages of both cluster and field stars of all ages and metallicities in the future.
@@ Acknowledgements.
Appendix A More [C/N] discussion
@@
Appendix B Chemical age estimates of NGC 6633

### v2

@@ 8 Summary, conclusions, and outlook
The large pixels will also reduce the quality of detailed asteroseismic studies in the crowded fields of clusters from the upcoming PLATO mission (Rauer et al., 2014), which is also not planned to observe the field of NGC 6633. Instead, future missions like STEP333 https://space.au.dk/the-space-research-hub/step/ and Haydn (Miglio et al., 2021b) will allow an extension of asteroseismic studies of cluster stars to a larger scale to ensure proper stellar ages of both cluster and field stars of all ages and metallicities in the future.
@@ Acknowledgements.
Appendix A Information on giant members of NGC 6633
Table 2: Information on giant members of NGC 6633
ID HD 170053 HD 170174 HD 170231 HD 170292 TIC 319770092
Symbol in plots cross square circle triangle diamond
Gaia DR3 ID 447746039- 447722337- 447727326- 447724911- 447725330-
1998886144 8527434880 8868842752 3972647168 5860760960
+0700329 +0636003 +0654514 +0642293 +0646000
Dec (degrees) 7.009091 6.600086 6.914277 6.708119 6.766667
[…]
$\nu$_eff 1.4085894 1.4590598 1.4565996 1.4639846 1.4669659
ecl_lat (deg) 30.263 29.84 30.158 29.947 30.007
RUWE 0.962 0.958 1.013 0.879 1.035
Parallax corr. (mas) -0.0329 -0.0328 -0.0326 -0.0325 -0.0324
RV ($\rm km\,s^{-1}$) -29.27(13) -28.71(13) -28.45(13) -29.02(14) -28.68(16)
$G$ (mag) 6.840 7.987 8.328 8.464 8.641
$K_{S}$ (mag)aaaaaaSkrutskie et al. (2006). 4.087 5.674 6.065 6.246 6.443
$G_{\rm BP}-G_{\rm RP}$ 1.567 1.295 1.293 1.259 1.242
E(BV)E(B-V) Green bbbbbbFrom Bayestar2019 E(gr)E(g-r), using the conversion factor 0.884 from Green et al. (2019), $A_{G}=2.74\times E(B-V)$ (Casagrande & VandenBerg 2018), $A_{K_{S}}=0.366\times E(B-V)$, $A_{G_{\rm BP}}=3.374\times E(B-V)$ and $A_{G_{\rm RP}}=2.035\times E(B-V)$ (Casagrande & VandenBerg 2014). 0.088 0.159 0.150 0.133 0.150
E(BV)E(B-V) Peña ccccccPeña et al. (2017). 0.182 0.182 0.182 0.182 0.182
E(BV)E(B-V) calc. $(V-K_{S})$ddddddAssuming adopted $T_{\rm eff}$s and requiring agreement with photometric temperatures from Casagrande & VandenBerg (2018). The one using $V-K_{S}$ is adopted. 0.048 0.180 0.184 0.177 0.180
E(BV)E(B-V) calc. $(G_{\rm BP}-G_{\rm RP})$ddddddAssuming adopted $T_{\rm eff}$s and requiring agreement with photometric temperatures from Casagrande & VandenBerg (2018). The one using $V-K_{S}$ is adopted. 0.112 0.185 0.205 0.189 0.187
$T_{\rm eff}$ (K)TsTsTsTsTsTsTsantaki et al. (2023). - 4924 5015 5068 5093
$T_{\rm eff}$ (K)MaMaMaMaMaMaMagrini et al. (2021). - 4957 5025 5056 -
$T_{\rm eff}$ (K)CaCaCaCaCaCaCasamiquela et al. (2021), numbers provided by L. Casamiquela (private comm.) - 5010 5095 5135 5165
$T_{\rm eff}$ (K)AdAdAdAdAdAdAdopted $T_{\rm eff}$s. See text. 4290 $50$ 5125 5160 5195
$T_{\rm eff}$ from $G_{\rm BP}-G_{\rm RP}$ (K)eeeeeeCalculated using Casagrande & VandenBerg (2018) assuming E(BV)E(B-V) from Green et al. (2019). [Fe/H] =+0.0=+0.0 and loggg = 2.7 was assumed. A change of $\pm 0.1$ dex in [Fe/H] corresponds to $\mp$8 K, a change of $\pm 0.5$ in loggg yields $\mp 2$ K 4240 4953 4933 4970 5110
$T_{\rm eff}$ from $G-K_{S}$ (K)eeeeeeCalculated using Casagrande & VandenBerg (2018) assuming E(BV)E(B-V) from Green et al. (2019). [Fe/H] =+0.0=+0.0 and loggg = 2.7 was assumed. A change of $\pm 0.1$ dex in [Fe/H] corresponds to $\mp$8 K, a change of $\pm 0.5$ in loggg yields $\mp 2$ K 4368 4987 5023 5029 5104
BCG (mag)ffffffAssuming the adopted $T_{\rm eff}$s, corresponding E(BV)E(B-V)s, Casagrande & VandenBerg (2018), and Gaia Parallaxes. -0.2603 0.0045 0.0194 0.0259 0.0321
[…]
$L(L_{\odot})$ from $K_{S}$ffffffAssuming the adopted $T_{\rm eff}$s, corresponding E(BV)E(B-V)s, Casagrande & VandenBerg (2018), and Gaia Parallaxes. 321.3(53) 122.4(21) 84.7(13) 75.0(12) 63.35(93)
$R_{SB}(R_{\odot})$ffffffAssuming the adopted $T_{\rm eff}$s, corresponding E(BV)E(B-V)s, Casagrande & VandenBerg (2018), and Gaia Parallaxes. 32.5(13) 14.46(48) 11.68(38) 10.85(35) 9.83(32)
A(Li) NLTEMoMaTsMoMaTsMoMaTsMoMaTsMoMaTsMoMaTsA(Li) NLTE from Table A5 of Morel et al. (2014) who used NLTE corrections from Lind et al. (2009). For the measurements marked with a *, the abundance was calculated using the numbers of Magrini et al. (2021) or Tsantaki et al. (2023) and the mean difference for the stars in common. 1.44(12) 0.80(12) 1.49(12) 1.21* 1.42*
A(Li) LTEMaMoTsMaMoTsMaMoTsMaMoTsMaMoTsMaMoTsA(Li) LTE from Magrini et al. (2021). For the measurements marked with a *, the abundance was calculated using the numbers of Morel et al. (2014); Lagarde et al. (2015) or Tsantaki et al. (2023) and the mean difference for the stars in common. 1.17* 0.66(5) 1.31(3) 1.05(3) 1.28*
A(Li) NLTEMaMoWaTsMaMoWaTsMaMoWaTsMaMoWaTsMaMoWaTsMaMoWaTsA(Li) 3D NLTE caluclated from Magrini et al. (2021) LTE values using 3D NLTE corrections from Wang et al. (2021). For the measurements marked with a *, the 3D NLTE correction was applied to the LTE value of Morel et al. (2014); Lagarde et al. (2015) or Tsantaki et al. (2023) corrected to Magrini et al. (2021) using the mean difference for the stars in common. We adopt the larger of the error-bars from Morel et al. (2014) that includes systematic errors. 0.97(12)* 0.61(12) 1.26(12) 1.00(12) 1.23(12)*
444 \$MoLaMa\$\$MoLaMa\$footnotetext: A(Li) LTE from the online Table A5 of Morel et al. (2014). For the measurements marked with a *, the abundance was calculated using the numbers from Magrini et al. (2021) and the mean difference for the stars in common. \$Ts\$\$Ts\$footnotetext: A(Li) LTE from Tsantaki et al. (2023)
Appendix B More [C/N] discussion
@@
Appendix C Chemical age estimates of NGC 6633